Three years ago, Marissa Moore started a business that creates and delivers holiday and birthday gift baskets to students at the local university. Marissa sells the baskets for $25 each, and her variable costs are $15 per basket. She incurs $13,800 in fixed costs each year. Last year, Marissa sold 4,020 baskets, and she believes that demand this year will be stable at 4,020 baskets. The following are the actions Marissa could take if she wants to earn $28,400 in operating income by selling only 4,020 baskets. Consider each action independently. (Round per unit answers to 2 decimal places, e.g. 52.75 and fixed cost to 0 decimal places, e.g. 5,275.)

1. Raise selling price per unit to $

2. Reduce variable costs per unit to $

3. Reduce fixed costs to $
Answer #1

1. Selling price to earn $28,400 operating income
Required contribution to earn desired operating income = Desired operating income + Fixed costs = $28,400 + $13,800 = $42,200
Variable costs for 4,020 baskets = Variable costs per unit * 4,020 units = $15 per unit * 4,020 units = $60,300
Required sales revenue for desired operating income = Required contribution margin + Variable costs = $60,300 + $42,200 = $102,500
Required Selling price per unit = Required sales revenue for desired operating income / 4,020 units = $102,500 / 4,020 units = $25.50
Increase in selling price = Required selling price - Current selling price = $25.50 - $25 = $0.50 per unit.

2. Variable costs per unit to earn $28,400 operating income
Required contribution to earn desired operating income = Desired operating income + Fixed costs = $28,400 + $13,800 = $42,200
Current sales revenue = 4,020 baskets * Current selling price = 4,020 * $25 per unit = $100,500
Total variable costs which will result in required operating income = Current sales revenue - Required contribution = $100,500 - $42,200 = $58,300
Variable costs per unit = Total variable costs / 4,020 units = $58,300 / 4,020 = $14.50 per unit.
Variable costs to be reduced to $14.50 to achieve desired operating income.

3. Fixed costs to earn $28,400 operating income
Current contribution margin = ( Selling price - Variable costs ) * 4,020 units = ( $25 - $15 ) * 4,020 = $40,200
Desired operating income = $28,400
Required fixed costs = Current contribution margin - Desired operating income = $40,200 - $28,400 = $11,800
Fixed costs to be reduced to $11,800 to earn $28,400 operating income.